OpenStudy (anonymous):

can anyone help i did it and i did not get the answer. Solve x2 – 6x = 27 by completing the square.

OpenStudy (anonymous):

please shoe work so i can do the next ones

OpenStudy (anonymous):

x^2 -6x-27 =0 x^2 -6x-9 -18=0 (x-3)^2 -18 =0

OpenStudy (anonymous):

(x-3)^2 =18 \[x-3=+- \sqrt{18} = +-\sqrt{9}\sqrt{2}\]

OpenStudy (anonymous):

\[x= 3 +- 3\sqrt{2}\]

OpenStudy (anonymous):

what are the two last post u did

OpenStudy (anonymous):

both side add 9: \[x^2-6x+9=27+9\] then the left side will become: \[x^2-6x+9=(x-3)^2\] then you got the answer: \[(x-3)^2=36\] or subtract 36 from both side: \[(x-3)^2-36=0\]